# Joint Value Estimation and Bidding in Repeated First-Price Auctions

### v1

@@ Appendix A Proofs of Main Lemmas
A.6 Proof of Lemma 5.3
Lemma A.7 (Restatement of Lemma 5.3).
Suppose the event in Lemma 5.2 holds. Let $\widehat{r}_{t,0}$ and $\widehat{r}_{t,1}$ be defined as in (14) and (15). For any $s\in[S]$ and for the criterion $i\in\{0,1\}$ selected in stage $s$ in Algorithm 5, it holds that
@@
where $b^{(s)}_{*,i}=\operatorname*{arg\,max}_{b\in B_{s}}u^{(s)}_{t,i}(b)$ for $i=0,1$ as defined in Algorithm 4.
@@
Fix any $s\in[S]$. For the ease of notation, we suppress the superscript $(s)$ for $A_{t}^{(s)}$, $b^{(s)}_{*,i}$, $w_{t}^{(s)}$, $u^{(s)}_{t,1}$, and $u^{(s)}_{t,0}$ in the remaining proof. Denote $\gamma_{t}=1+c\log(2T)+c^{\prime}\sqrt{\sum_{\tau\in\Phi_{t}^{(s)}}\delta_{\tau}^{2}}$ as defined in Algorithm 4. By Lemma C.1, when $2\gamma_{t}\|x_{t}\|_{A_{t}^{-1}}\leq\frac{1}{20L}$, we have $\widehat{G}_{t}(b_{*,0})-\widehat{G}_{t}(b_{*,1})\leq 1-\frac{1}{20L}$.
@@
A.7 Proof of Lemma 5.5
Lemma A.8 (Restatement of Lemma 5.5).
In Algorithm 5, for every $t\in[T]$ and $s\in[S]$, it holds that $\max_{b^{*}\in{\mathcal{B}}}\widehat{r}_{t}(b^{*})-\widehat{r}_{t}(b)\leq 8\cdot 2^{-s}$ for all $b\in B_{s}$.
@@
Suppose the hypothesis holds for $s-1$. The elimination step (Line 3) in Algorithm 3 implies that for any $b\in B_{s}$, we have $w^{(s-1)}_{t}(b)\leq 2^{-(s-1)}$, $w^{(s-1)}_{t}(b^{*}_{t})\leq 2^{-(s-1)}$, and by Lemma 5.3 $u_{t,i}^{(s-1)}(b)\geq u_{t,i}^{(s-1)}(b^{*}_{t})-2^{-s+2}$ for the criterion $i\in\{0,1\}$ selected in stage $s-1$. Since $b^{*}_{t}\in B_{s-1}$ by the induction hypothesis, by Lemma 5.3,
@@
where (a) uses Lemma 4.1 and (b) is by the elimination criterion in Algorithm 5. This concludes the induction and hence the proof. ∎
A.8 Proof of Lemma 5.6
Then we are ready for the proof of Lemma 5.6.
Lemma A.9 (Restatement of Lemma 5.6).
@@
for an absolute constant $c>0$. Note that Assumption 2.7 implies $\|\widehat{G}_{t}-G_{t}\|_{\infty}\leq 2\delta_{t}$. Thus by the truncation step in Line 22 of Algorithm 5 and Lemma C.4, it holds that
@@
since $\delta_{t}<0.1$; and (b) uses Lemma A.3 and that $\gamma_{t}\leq\gamma$. ∎
A.9 Proof of Lemma 6.2
The proof of Lemma 6.2 is a direct consequence of Lemma A.3 given the following explicit expression of $\{\delta_{t}\}$:
@@
Under the conditions of Lemma 6.2, there is an estimator $\widehat{G}_{t}$ for $G_{t}$ that satisfies Assumption 2.7 with
@@
Given the Lipschitzness in Assumption 2.3, we consider the discretized bid space ^=$\widehat{{\mathcal{B}}}=\{\frac{i}{T}:i\in\{0\}\cup[T]\}$. At each time $t$, we estimate the parameter $\varphi_{*}$ via linear regression with a careful data splitting scheme: Denote $\Sigma_{t}=18I+\sum_{s<t}x_{s}x_{s}^{\top}$. By Lemma C.7, there is a subset $S_{t}\subseteq[t-1]$ such that $|S_{t}|\leq\frac{t-1}{2}$ and $18I+\sum_{s\in S_{t}}x_{s}x_{s}^{\top}\succeq\frac{\Sigma_{t}}{9}$. Define $A_{t}=18I+\sum_{s\in S_{t}}x_{s}x_{s}^{\top}$ and $z_{t}=\sum_{s\in S_{t}}M_{s}x_{s}$, and set the estimator $\widehat{\varphi}_{t}=A_{t}^{-1}z_{t}$. By Lemma 4.1, with probability at least $21-T^{-2}/$ 2, we have a confidence bound that
@@
for $\beta=O(\sigma\sqrt{\log T})$ and all $x\in\{x_{t}-x_{s}:s\in[t-1]\}$, where $\sigma$ is the sub-Gaussian parameter in Assumption 6.1. Construct the empirical CDF as follows: let $X_{s}(b)=\mathds{1}[M_{s}-\widehat{\varphi}_{t}^{\top}x_{s}+\widehat{\varphi}_{t}^{\top}x_{t}\leq b]$ and $\widehat{G}_{t}(b)=\frac{1}{|S_{t}^{c}|}\sum_{s\in S_{t}^{c}}X_{s}(b)$ using the remaining data $S_{t}^{c}=[t-1]\backslash S_{t}$. Note
@@
Write the target value as $v=\varphi_{*}^{\top}x_{t}-b$ for simplicity. Since ${\mathbb{E}}[X_{s}(b)]=Q(v+\varepsilon_{s})$, its variance is $\mathsf{Var}(X_{s}(b))=Q(v+\varepsilon_{s})(1-Q(v+\varepsilon_{s}))$. Now we apply Bernstein’s inequality (cf. Lemma C.5) to $\widehat{G}_{t}(b)$ and a union bound to obtain, with probability at least $21-T^{-2}/$ 2, for all $b\in\widehat{{\mathcal{B}}}$,
@@
Let $a_{0},a_{1},a_{2}>0$ be the constants in Assumption 6.1. When $Q(v)\in(a_{0},1-a_{0})$, we bound the second term as
@@
for a constant $c>0$, where (a) uses Lemma A.3 and the properties of $S_{t}$. On the other hand, when $Q(v)\in[0,a_{0}]\cup[1-a_{0},a]$, we have
@@
Plugging (29) and (30) into (28) and substituting in $|S_{t}^{c}|\geq\frac{t-1}{2}\geq\frac{t}{4}$, the CDF estimation error is bounded as desired: for every $b\in\widehat{{\mathcal{B}}}$, with probability at least $21-T^{-2}/$ 2,
@@
For any $b\in{\mathcal{B}}$ in the continuous bid space, let its discretized neighbors be $h_{1}=h_{2}-T^{-1}\leq b\leq h_{2}$ for $h_{1},h_{2}\in\widehat{{\mathcal{B}}}$. Thanks to the Lipschitzness of $Q$, for both $h=h_{1},h_{2}$, it holds that $G_{t}(h)(1-G_{t}(h))\leq G_{t}(b)(1-G_{t}(b))+\frac{L}{T}$. Hence by monotonicity of $\widehat{G}_{t}$ and the triangular inequality, with probability at least $21-T^{-2}/$ 2,
@@ Appendix B Proof of lower bounds
In this section we prove the regret lower bounds in Theorem 2.9. We begin with the analysis of Model 1.
@@
(32)${\geq}\frac{\Delta}{2}\sum_{t=1}^{T}\int\min\left\{\mathrm{d}\mathbb{P}_{1}^{\otimes t},\mathrm{d}\mathbb{P}_{2}^{\otimes t}\right\}$
=$\overset{\rm(a)}{=}\frac{\Delta}{2}\sum_{t=1}^{T}\left(1-\|\mathbb{P}_{1}^{\otimes t}-\mathbb{P}_{2}^{\otimes t}\|_{\mathrm{TV}}\right)$
@@
Note that Theorem B.1 completes the proof of Theorem 2.9 for Model 1, with $L=O(1)$ in Assumption 2.3. We proceed to show that the special lower bound instance in Theorem B.1 can also be embedded in other lower bounds of Theorem 2.9. Under Model 2 and $d\geq 2$, let
@@
with $\Delta=\frac{1}{4}\sqrt{\frac{d-1}{T}}$. Since $T\geq d^{2}$, it holds that $\|\theta_{1,*}\|_{2}\leq 1$ almost surely. Divide the time horizon $T$ into $d-1$ sub-horizons $T_{i}$ for $i=1,\dots,d-1$ with equal length $\frac{T}{d-1}$, and let
@@
during the sub-horizon $T_{i}$, where the second $\frac{1}{2}$ appears in the $(i+1)$-th entry. As for $M_{t}$, we again use the i.i.d. distribution in Theorem B.1. Consequently, the regret $\mathsf{Reg}_{\mathrm{lpo}}(\pi)$ can be decomposed as the sum of regrets from $d-1$ independent sub-problems, each of time duration $\frac{T}{d-1}$. By Theorem B.1, we have
@@
as claimed. For the case of $d=1$, we simply use a different construction $\theta_{1,*}\sim\mathrm{Unif}(\{\frac{1}{4}-\Delta,\frac{1}{4}+\Delta\})$ and $x_{t}\equiv 1$, so that Theorem B.1 again gives a regret lower bound $\Omega(\sqrt{T})$. Finally, since $\mathsf{Reg}_{\mathrm{lte}}(\pi)\geq\mathsf{Reg}_{\mathrm{lpo}}(\pi)$, the lower bound for Model 3 also follows.
@@ Appendix C Auxiliary lemmas
For simplicity, write $b_{1}=\widehat{b}_{*}(v_{1})$ and $b_{2}=\widehat{b}_{*}(v_{2})$. By Lemma C.2, we have $\widehat{G}(b_{1})\leq\widehat{G}(b_{2})$. For the sake of contradiction, assume $\widehat{G}(b_{2})-\widehat{G}(b_{1})>1-\frac{1}{20L}$, which implies $\widehat{G}(b_{1})\leq\frac{1}{20L}$ and $\widehat{G}(b_{2})>1-\frac{1}{20L}$.
@@
In Algorithm 5, at any time $t$ and any stage $s$, it holds that
@@
At each stage $s$, let $\hat{\theta}_{t}$ be the estimated parameter in Algorithm 4. By Lemma 5.2, we have $\left\lvert\hat{\theta}_{t}^{\top}x_{t}-\theta_{*}^{\top}x_{t}\right\rvert\leq\gamma\|x_{t}\|_{{A_{t}^{(s)}}^{-1}}$, where $\gamma$ is defined in Line 7 of Algorithm 4. Then since the mapping $v\mapsto\operatorname*{arg\,max}_{b\in{\mathcal{B}}}G(b)\left(v-b\right)$ is monotone by Lemma C.2, the claim follows from the definitions of $b^{(s)}_{*,0}$ and $b^{(s)}_{*,1}$ in Algorithm 4 and that $\widehat{r}_{t}(b^{*})=\widehat{G}_{t}(b^{*})(\theta_{*}^{\top}x_{t}-b^{*})$. ∎

### v2

@@ Appendix A Proofs of Main Lemmas
A.6 Proof of Lemma˜5.3
Lemma A.7 (Restatement of Lemma˜5.3).
Suppose the event in Lemma 5.2 holds. Let $\widehat{r}_{t,0}$ and $\widehat{r}_{t,1}$ be defined as in (14) and (15). For any $s\in[S]$ and for the criterion $i\in\{0,1\}$ selected in stage $s$ in Algorithm˜5, it holds that
@@
where $b^{(s)}_{*,i}=\operatorname*{arg\,max}_{b\in B_{s}}u^{(s)}_{t,i}(b)$ for $i=0,1$ as defined in Algorithm˜4.
@@
Fix any $s\in[S]$. For the ease of notation, we suppress the superscript $(s)$ for $A_{t}^{(s)}$, $b^{(s)}_{*,i}$, $w_{t}^{(s)}$, $u^{(s)}_{t,1}$, and $u^{(s)}_{t,0}$ in the remaining proof. Denote $\gamma_{t}=1+c\log(2T)+c^{\prime}\sqrt{\sum_{\tau\in\Phi_{t}^{(s)}}\delta_{\tau}^{2}}$ as defined in Algorithm˜4. By Lemma C.1, when $2\gamma_{t}\|x_{t}\|_{A_{t}^{-1}}\leq\frac{1}{20L}$, we have $\widehat{G}_{t}(b_{*,0})-\widehat{G}_{t}(b_{*,1})\leq 1-\frac{1}{20L}$.
@@
A.7 Proof of Lemma˜5.5
Lemma A.8 (Restatement of Lemma˜5.5).
In Algorithm˜5, for every $t\in[T]$ and $s\in[S]$, it holds that $\max_{b^{*}\in{\mathcal{B}}}\widehat{r}_{t}(b^{*})-\widehat{r}_{t}(b)\leq 8\cdot 2^{-s}$ for all $b\in B_{s}$.
@@
Suppose the hypothesis holds for s1s-1. The elimination step (Line 3) in Algorithm˜3 implies that for any $b\in B_{s}$, we have $w^{(s-1)}_{t}(b)\leq 2^{-(s-1)}$, $w^{(s-1)}_{t}(b^{*}_{t})\leq 2^{-(s-1)}$, and by Lemma 5.3 $u_{t,i}^{(s-1)}(b)\geq u_{t,i}^{(s-1)}(b^{*}_{t})-2^{-s+2}$ for the criterion $i\in\{0,1\}$ selected in stage s1s-1. Since $b^{*}_{t}\in B_{s-1}$ by the induction hypothesis, by Lemma˜5.3,
@@
where (a) uses Lemma˜4.1 and (b) is by the elimination criterion in Algorithm˜5. This concludes the induction and hence the proof. ∎
A.8 Proof of Lemma˜5.6
Then we are ready for the proof of Lemma˜5.6.
Lemma A.9 (Restatement of Lemma˜5.6).
@@
for an absolute constant $c>0$. Note that Assumption˜2.7 implies $\|\widehat{G}_{t}-G_{t}\|_{\infty}\leq 2\delta_{t}$. Thus by the truncation step in Line 22 of Algorithm 5 and Lemma˜C.4, it holds that
@@
since $\delta_{t}<0.1$; and (b) uses Lemma˜A.3 and that $\gamma_{t}\leq\gamma$. ∎
A.9 Proof of Lemma˜6.2
The proof of Lemma˜6.2 is a direct consequence of Lemma˜A.3 given the following explicit expression of $\{\delta_{t}\}$:
@@
Under the conditions of Lemma˜6.2, there is an estimator $\widehat{G}_{t}$ for $G_{t}$ that satisfies Assumption˜2.7 with
@@
Given the Lipschitzness in Assumption˜2.3, we consider the discretized bid space $\widehat{{\mathcal{B}}}=\{\frac{i}{T}:i\in\{0\}\cup[T]\}$. At each time $t$, we estimate the parameter $\varphi_{*}$ via linear regression with a careful data splitting scheme: Denote $\Sigma_{t}=18I+\sum_{s<t}x_{s}x_{s}^{\top}$. By Lemma˜C.7, there is a subset $S_{t}\subseteq[t-1]$ such that $|S_{t}|\leq\frac{t-1}{2}$ and $18I+\sum_{s\in S_{t}}x_{s}x_{s}^{\top}\succeq\frac{\Sigma_{t}}{9}$. Define $A_{t}=18I+\sum_{s\in S_{t}}x_{s}x_{s}^{\top}$ and $z_{t}=\sum_{s\in S_{t}}M_{s}x_{s}$, and set the estimator $\widehat{\varphi}_{t}=A_{t}^{-1}z_{t}$. By Lemma˜4.1, with probability at least $1-T^{-2}/2$, we have a confidence bound that
@@
[…]
@@
Write the target value as $v=\varphi_{*}^{\top}x_{t}-b$ for simplicity. Since ${\mathbb{E}}[X_{s}(b)]=Q(v+\varepsilon_{s})$, its variance is $\mathsf{Var}(X_{s}(b))=Q(v+\varepsilon_{s})(1-Q(v+\varepsilon_{s}))$. Now we apply Bernstein’s inequality (cf. Lemma C.5) to $\widehat{G}_{t}(b)$ and a union bound to obtain, with probability at least $1-T^{-2}/2$, for all $b\in\widehat{{\mathcal{B}}}$,
@@
Let $a_{0},a_{1},a_{2}>0$ be the constants in Assumption˜6.1. When $Q(v)\in(a_{0},1-a_{0})$, we bound the second term as
@@
for a constant $c>0$, where (a) uses Lemma˜A.3 and the properties of $S_{t}$. On the other hand, when $Q(v)\in[0,a_{0}]\cup[1-a_{0},a]$, we have
@@
Plugging (29) and (30) into (28) and substituting in $|S_{t}^{c}|\geq\frac{t-1}{2}\geq\frac{t}{4}$, the CDF estimation error is bounded as desired: for every $b\in\widehat{{\mathcal{B}}}$, with probability at least $1-T^{-2}/2$,
@@
For any $b\in{\mathcal{B}}$ in the continuous bid space, let its discretized neighbors be $h_{1}=h_{2}-T^{-1}\leq b\leq h_{2}$ for $h_{1},h_{2}\in\widehat{{\mathcal{B}}}$. Thanks to the Lipschitzness of $Q$, for both $h=h_{1},h_{2}$, it holds that $G_{t}(h)(1-G_{t}(h))\leq G_{t}(b)(1-G_{t}(b))+\frac{L}{T}$. Hence by monotonicity of $\widehat{G}_{t}$ and the triangular inequality, with probability at least $1-T^{-2}/2$,
@@ Appendix B Proof of lower bounds
In this section we prove the regret lower bounds in Theorem˜2.9. We begin with the analysis of Model 1.
@@
[…]
$\overset{\rm(a)}{=}\frac{\Delta}{2}\sum_{t=1}^{T}\left(1-\|\mathbb{P}_{1}^{\otimes t}-\mathbb{P}_{2}^{\otimes t}\|_{\mathrm{TV}}\right)$
@@
Note that Theorem˜B.1 completes the proof of Theorem˜2.9 for Model 1, with $L=O(1)$ in Assumption˜2.3. We proceed to show that the special lower bound instance in Theorem˜B.1 can also be embedded in other lower bounds of Theorem˜2.9. Under Model 2 and $d\geq 2$, let
@@
with $\Delta=\frac{1}{4}\sqrt{\frac{d-1}{T}}$. Since $T\geq d^{2}$, it holds that $\|\theta_{1,*}\|_{2}\leq 1$ almost surely. Divide the time horizon $T$ into d1d-1 sub-horizons $T_{i}$ for $i=1,\dots,d-1$ with equal length $\frac{T}{d-1}$, and let
@@
during the sub-horizon $T_{i}$, where the second $\frac{1}{2}$ appears in the $(i+1)$-th entry. As for $M_{t}$, we again use the i.i.d. distribution in Theorem˜B.1. Consequently, the regret $\mathsf{Reg}_{\mathrm{lpo}}(\pi)$ can be decomposed as the sum of regrets from d1d-1 independent sub-problems, each of time duration $\frac{T}{d-1}$. By Theorem˜B.1, we have
@@
as claimed. For the case of $d=1$, we simply use a different construction $\theta_{1,*}\sim\mathrm{Unif}(\{\frac{1}{4}-\Delta,\frac{1}{4}+\Delta\})$ and $x_{t}\equiv 1$, so that Theorem˜B.1 again gives a regret lower bound $\Omega(\sqrt{T})$. Finally, since $\mathsf{Reg}_{\mathrm{lte}}(\pi)\geq\mathsf{Reg}_{\mathrm{lpo}}(\pi)$, the lower bound for Model 3 also follows.
@@ Appendix C Auxiliary lemmas
For simplicity, write $b_{1}=\widehat{b}_{*}(v_{1})$ and $b_{2}=\widehat{b}_{*}(v_{2})$. By Lemma˜C.2, we have $\widehat{G}(b_{1})\leq\widehat{G}(b_{2})$. For the sake of contradiction, assume $\widehat{G}(b_{2})-\widehat{G}(b_{1})>1-\frac{1}{20L}$, which implies $\widehat{G}(b_{1})\leq\frac{1}{20L}$ and $\widehat{G}(b_{2})>1-\frac{1}{20L}$.
@@
In Algorithm˜5, at any time $t$ and any stage $s$, it holds that
@@
At each stage $s$, let $\hat{\theta}_{t}$ be the estimated parameter in Algorithm˜4. By Lemma 5.2, we have $\left\lvert\hat{\theta}_{t}^{\top}x_{t}-\theta_{*}^{\top}x_{t}\right\rvert\leq\gamma\|x_{t}\|_{{A_{t}^{(s)}}^{-1}}$, where $\gamma$ is defined in Line 7 of Algorithm˜4. Then since the mapping $v\mapsto\operatorname*{arg\,max}_{b\in{\mathcal{B}}}G(b)\left(v-b\right)$ is monotone by Lemma C.2, the claim follows from the definitions of $b^{(s)}_{*,0}$ and $b^{(s)}_{*,1}$ in Algorithm 4 and that $\widehat{r}_{t}(b^{*})=\widehat{G}_{t}(b^{*})(\theta_{*}^{\top}x_{t}-b^{*})$. ∎